Davion 《H01 C) 2 of 1
Answer #1

a) on increasing volume of the crate, the mass also increases hence the density of the material remained unchanged,

ratio = 1

=======

b)

as we know  

density = mass / volume

for mass = constant

density is inversely proportional to volume

D' / D = V / V'

D' / D = L^3 / ( 2L)^3

D'/ D = 1/8

=======

do comment in case any doubt will reply will reply for sure.. Goodluck